OpenStudy (anonymous):

Which of the following equations represents an inverse variation? y equals x divided by two y equals two times x y equals two divided by x y equals x squared

OpenStudy (radar):

Which one would cause y to decrease as x increased?

OpenStudy (anonymous):

umm dividing the number by x

OpenStudy (radar):

Yes y=2/x when x = 1 y would be 2 now increase x to 2 then y=2/2 = 1 it is decreasing. 3rd choice is the way to go.

OpenStudy (anonymous):

ok so it's y = 2/x right

OpenStudy (radar):

absolutely, as the x get bigger, the y will get smaller ("inverse")

OpenStudy (radar):

All the rest as x gets bigger y also get bigger, vice-versa

OpenStudy (anonymous):

so u mean switch it around like x/2

OpenStudy (radar):

No NO NOO, 2/x is an INVERSE variation which means as x gets bigger y gets smaller or as x gets smaller y gets bigger. y=x/2 is a DIRECT variation and as x gets bigger y will get bigger, as x gets smaller, y will get smaller.

OpenStudy (anonymous):

Oh I get it now , my meta-math was messing me up so the answer is A right

OpenStudy (radar):

No.

OpenStudy (anonymous):

I really don't understand why it not

OpenStudy (radar):

Here is why, A is y= x/2 Now let us test this: If x = 2 what will y be?

OpenStudy (radar):

Earth to Lyssa123 What will y be when x=2 in A.

OpenStudy (radar):

y = 2/2 = 1 Do you follow that?

OpenStudy (anonymous):

Kinda

OpenStudy (radar):

Now lets make x bigger, lets make it 4 (4 is bigger than 2) y=4/2=2 Don't you see that y got bigger?, it went from a 1 to a 2. this is not rocket science, it is simple arithmetic.

OpenStudy (radar):

So as x gets bigger so does y, this is not a inverse variation it is a direct variation.

OpenStudy (radar):

That is why A or the first choice is a bad one. Please start from my first post and please read them and attempt to follow.

OpenStudy (radar):

I have already told you the answer in a previous post, trust me.

OpenStudy (anonymous):

Ok see what u is saying , i just need to read slowly and C is the right answer is what u is try to say

OpenStudy (radar):

Yes, but you see I can't say C, cause then I would be giving away the answer.

OpenStudy (anonymous):

ok

OpenStudy (radar):

That would be a violation of the OS rules. Besides I don't see any C or A or B or D.

OpenStudy (radar):

Good luck with your studies. Oh by the way I use to live in St. Simons Island, Ga........Not too far from Jacksonville,FL

OpenStudy (anonymous):

Thanks and Kool